100- Relative Intensity T 10 15 20 25 30 35 50 55 60 65 70 75 40 45 m/z
Answer #1

IR stretching around 1715 cm-1 indicate the presence of a carbonyl group from the aliphatic ketone.

The peak near 2950 cm-1 is due to C-H stretching frequency.

Absence of peak between 1500-1600 indicates that the C=C group is not present in the molecule.

From the Mass spectra,

m/z = 43 is could be because of phpLxJLJ9.png fragment.

The m/z = 57 could be because of 06 fragment.

From the above information and Molecular ion peak at m/z = 72 indicate the compound  phppK4mWe.png